OpenStudy (korosh23):

Calculus Question (Limits and Continuity) I need to understand why this rational function is continuous ?

OpenStudy (korosh23):

In this question we can observe that in the denominator, there is a restriction and that is x cannot equal 5/3. Ok, now we can understand that there is a vertical asymptote at x = 5/3. Question: Does this mean that this rational function in the form of f(x) = P(x) / Q(x) is continuous ?

OpenStudy (korosh23):

According to a theorem in my textbook, it says that all rational functions are continuous functions, but we just saw that there is a vertical asymptote at x =5/3, and that creates discontinuity. Am I right ?

Nnesha (nnesha):

yes, the function is continuous for all x values except where denominator equation equals zero

Nnesha (nnesha):

i'm sorry plz ignore the word equation

OpenStudy (korosh23):

ok, so you mean there is a disconinuity when the denominaotr equals zero. Doesn't that make the whole function discontinous? Because now the function is not connected and is not united as a whole. Like there is a cut in between a line.

Nnesha (nnesha):

yes the function `f` would be discontinuous at x=3/5

Nnesha (nnesha):

numerator is continuous x^3+2x^2-1 and denominator is continuous 5-3x and f(x)= (x^3+2x^2-1)/( 5-3x) is continuous for all x values except x=3/5

OpenStudy (korosh23):

Ok, so you mean the point of discontinuity at x = 5/3 will not affect the whole continuity of the function. Got it :)

OpenStudy (korosh23):

Thanks for your help

Nnesha (nnesha):

*whole continuity of the function * do you mean f(x) function right ? x^3+2x^2-1 divided by 5-3x ?

OpenStudy (korosh23):

Yes, exactly. In my opinion, the vertical asymptote will not affect the whole continuity, it will only make the function f(x) discontinuous at a specific point x = 5/3.

Nnesha (nnesha):

okay now make sense i was confused bec you said ` whole function discontinous` yes right at that point function is discontinous

Nnesha (nnesha):

i thought you were trying to ask at x=3/5 *whole function* discontinuous or just denominator

OpenStudy (korosh23):

I was trying to see if with the point of discontinuity at x = 5/3, would the function as a whole stay continuous which is yes. I understood the concept now. Your explanation was helpful.:)

Nnesha (nnesha):

:=))

OpenStudy (danjs):

"According to a theorem in my textbook, it says that all rational functions are continuous functions" Look at the rule again there is a part missing from yours...

korosh23:

I asked this question back in OS. lol :D

Ultrilliam:

Yes, yes you did

KjSaif:

????

SkyVoltage43:

why is everyones name who post a question OpenStudy?

korosh23:

OpenStudy user is created to upload all the questions asked back in OpenStudy, so people can have once again access to it.

korosh23:

However, only the answers of the user who asked the question is posted and not the person who answered the question.

SkyVoltage43:

ok